OpenStudy (anonymous):

If 3t - s _________ 4 What is the value of s in terms of t

OpenStudy (anonymous):

Oh and the expressionequals 8s

OpenStudy (hwyl):

WHAT IS THAT? CAN YOU TYPE THE PROBLEM PROPERLY

OpenStudy (anonymous):

Its 3t-s/4 =8s @hwyl

OpenStudy (texaschic101):

(3t - s) / 4 = 8s --- multiply both sides by 4 3t - s = 32s -- add s to both sides 3t = 33s -- divide both sides by 3 t = 11s

OpenStudy (texaschic101):

or t = 11s s = t/11

OpenStudy (anonymous):

Thank youuu

OpenStudy (texaschic101):

YW :)